OpenStudy (anonymous):

Medal & Fan!(: Which equation can be used to solve the following word problem? Emilio has 6 more nickels than dimes, and the total value of his coins is $2.10. How many nickels does he have? Let n represent the number of nickels. 0.10n + 0.05n + 6 = 2.10 0.10n + 0.05(n + 6) = 2.10 0.10(n + 6) + 0.05n = 2.10 0.10(n – 6) + 0.05n = 2.10

OpenStudy (anonymous):

nickels = dimes + 6 n = d+6 d = n - 6

OpenStudy (amilapsn):

what's the value of a nickel/

OpenStudy (anonymous):

You have you have n-6 dimes, with that said, which one of your answers comes the closest to that...

OpenStudy (kamibug):

A nickel is 5 cents. (0.05) a dime is 10 cents (0.10). So those values will be their coefficients. (0.05)n + (0.10)d And we can set it equal to $2.10 because that's what the total of coins is worth. (0.05)n + (0.10)d = 2.10 And we can change d to n - 6 because he has 6 more nickels than dimes. (0.05)n + (0.10)(n - 6) = 2.10